Q9. (a) What is the potential between two points situated 10 cm and 20 cm from a 3.0 μC point charge?

(b)To what location should the point at 20 cm be moved to increase this potential difference by a factor of two?


1
Expert's answer
2021-01-10T18:31:19-0500

The potential difference between points from a point charge is expressed by the equation

"V_A -V_B = kQ(\frac{1}{r_a} - \frac{1}{r_b})"

Constant "k = 9 \times 10^9 \; N \; m^2/C^2"

ra is position of first point

rb is position of second point

(a)

"V_A -V_B = (9 \times 10^9 \;Nm^2/C^2)(3.0 \; μC)(\frac{1\;C}{10^6 \;μC})(\frac{1}{0.1 \;m} - \frac{1}{0.2 \;m}) \\

= 135 \times 10^3 \;V"

The potential between two points situated 10 cm and 20 cm from a 3.0 μC point charge is "135 \times 10^3 \;V" .

(b)

"\frac{1}{r_b}=\frac{1}{r_a}- \frac{V}{kQ} \\

\frac{1}{r_b}=\frac{1}{0.1 \;m}- \frac{2 \times 135 \times 10^3 \;V}{(9 \times 10^9 \;Nm^2/C^2)(3.0 \times 10^{-6}\;C)} \\

\frac{1}{r_b}= 0 \\

r_b = \infin"

Therefore, to double the potential between the points, the 20 cm point has to be shift to infinity.